OpenStudy (anonymous):

How woule I solve 2x^2 + 4x - 30 = 0 ?

OpenStudy (ddcamp):

(2x - a ) ( x + b) 2b - a = 4 ab = 30

OpenStudy (anonymous):

divide by 2 first gives x^2+2x-15=0 now get two factore that when multiplied you get 15 yet when added you get +2 5 and -3 use them to form (x+5)(x-3)=0 now seperate brackets x+5=0,gives x=-5 x-3=0 gives x=3